OpenStudy (goformit100):

Which of the following is closest to 27.8 × 9.6? A. 280 B. 300 C. 2,800 D. 3,000

OpenStudy (anonymous):

A) 280, because 27.8 x 9.6 = 266.88 and 280 is the closest.

OpenStudy (goformit100):

How to do it mentally ?

OpenStudy (goformit100):

Thank You

OpenStudy (anonymous):

\[ 27.8*9.6 \approx 28\times {9.5\times2 \over 2}=14\times19=2\times 7\times19=2\times 133=\mathbf{266} \]